# Understanding the Uncertainty of LLM Explanations: A Perspective Based on Reasoning Topology

### v1

@@ 5.3. Redundancy Measure of LLMs (RQ2)
Benefit from the topology structure, we are able to extract the reasoning path $\mathcal{P}_{\text{valid}}$ that successfully connects from nodeRaw to nodeResult. Then we can effectively compute the node that is not contributing to the final answer, and this serves as a sign of redundancy in the LLM’s reasoning process. Following the Eq. 21, we analyze the redundancy rate for each of the LLMs including GPT-4o-mini, Llama3-8b, Phi4, DeepSeek-R1(distilled), and Llama2-70b on both the node and edge redundancy. We find that, surprisingly, the GPT-4o-mini shows a significantly high redundancy rate in both the nodes and edges, it might reveal the high accuracy of the model comes from a border searching space (or generating length) when conducting reasoning and proposing solutions. However, this also reflects there would be a great potential to optimize the reasoning process for the GPT-wise models. Comparatively, the DeepSeek (distilled llama version) is relatively low, which might indicate the model’s training was conducted with a special design to encourage the ‘valid’ reasoning which eventually contributes to the final result. We show more results in the appendix for reference.
Figure 4. The quantification of the LLM’s reasoning redundancy on three datasets.
5.4. Findings on the Reasoning Patterns from LLMs (RQ3)
[…]
6. Conclusion
In this paper, we highlight the critical importance of understanding uncertainty in LLM explanations by revealing the reasoning consistency. By introducing a novel framework grounded in a reasoning topology perspective, the paper provides a structured method to decompose explanations into knowledge and reasoning dimensions. This allows for precise quantification of uncertainty, assessment of knowledge redundancy, and deeper insights into the model’s reasoning structure. Then, the paper based on the proposed formal structural construction, it propose a graph-edit distance based uncertainty measure, empirical study shows a better performance in revealing the true faithfulness of the natural language models. After that, a redundancy-based method is introduced to quantify the redundancy of the LLMs. The approaches not only enhance interpretability but also serve as a guide for improving robustness and faithfulness in LLM-generated explanations.
@@ Appendix A Notation Table
Table 3. The notations and explanations in this paper.
Appendix B Details of the GeoQA Dataset
@@
Appendix C Empirical Study on the Choice of Number for Generations
@@
Appendix D Details of Baseline methods
D.0.1. CoTA.
@@ Appendix D Details of Baseline methods
(22) $\text{CoTA}(CoT_{a},CoT_{b})=$ $\ \frac{1}{N_{a}+N_{b}}\bigg{(}\sum_{i=1}^{N_{a}}\max_{j=1,\dots,N_{b}}E(s_{ai},s_{bj})$
@@
Table 4. Examples of GeoQA dataset.
@@
Glacier These questions explore glacial movement, erosion, and the impact of climate change on ice dynamics.
@@
Striations indicate past movement direction, requiring cause-and-effect analysis, while glacier mass loss demands understanding the imbalance between accumulation and ablation.
@@
Earthquake These questions focus on seismic wave behavior, fault activity, and earthquake detection.
@@
Triangulation requires reasoning through wave arrival times, while P-wave detection relies on comparing wave speeds and impact to explain early warning systems.
@@
permafrost These questions examine permafrost thawing, climate feedback loops, and seasonal variations.
@@
Thawing permafrost and methane release involve feedback loops, while active layer variations require analyzing environmental factors like temperature and insulation.
@@
Monsoon These questions cover seasonal wind shifts, monsoon patterns, and storm formation.
@@
The East Asian monsoon’s impact on seasons involves reasoning about wind shifts, while monsoon depressions require linking low-pressure systems to storm formation.
@@
Figure 5. The model’s success rate to generate legitimate reasoning topology. It is calculated by the percentage of LLMs successfully generating the reasoning path from the nodeRaw to nodeResult following the few-shot prompt. It can be witnessed that, generally the GeoQA is a harder one to generate due to the conditional question types, which are rarely seen in the LLM training tasks.
Appendix E Research Methods
E.1. Part One
@@
(23) $\text{GED}(\mathcal{G}_{1},\mathcal{G}_{2})=C_{\text{sub.}}(\mathcal{P})+C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})$
@@
(24) $C_{\text{sub.}}(\mathcal{P})=\sum_{(v_{i},v_{j})\in\mathcal{P}_{v}}c(v_{i},v_{j})+\sum_{(e_{k},e_{m})\in\mathcal{P}_{e}}c(e_{k},e_{m}),$
where $c(v_{i},v_{j})$ and $c(e_{k},e_{m})$ represent the node and edge substitution costs, respectively. The term $C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})$ captures the deletion costs for nodes and edges in $\mathcal{G}_{1}$ that are not matched, given by:
(25) $C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})=\sum_{v_{i}\in\mathcal{V}_{1}\setminus\mathcal{P}_{v}}c_{\text{del.}}(v_{i})+\sum_{e_{k}\in\mathcal{E}_{1}\setminus\mathcal{P}_{e}}c_{\text{del.}}(e_{k}).$
Here, $c_{\text{del.}}(v_{i})$ and $c_{\text{del.}}(e_{k})$ denote the deletion costs of unmatched nodes and edges, respectively. This formulation quantifies the total cost required to align the two reasoning structures by summing the substitution costs of matched components and the deletion costs of unmatched ones.
E.2. Algorithm for Detection of Redundant Nodes and Dead Branches
@@
(26) $\text{Out}(v_{k})=\emptyset\quad\text{and}\quad v_{k}\neq\text{NodeResult}.$
@@
(27) $v_{k}\notin\bigcup_{[v_{i},v_{j},e_{ij}]\in\mathcal{P}_{\text{valid}}}\{v_{i},v_{j}\}.$
@@
Appendix F Experimental Details
F.1. Details of the dataset
@@
  • GeoQA: GeoQA is a self-constructed dataset designed to evaluate the reasoning capabilities of LLMs in conditional questions. With 20 categories, including climate and tectonic processes, its tasks require inference of specific results from given conditions. GeoQA emphasizes multi-step reasoning and domain-specific integration, making it a challenging benchmark (details in the Appendix B).
F.2. The calculation of the faithfulness core

### v2

@@ 5.3 Redundancy Measure of LLMs (RQ2)
Benefitting from the topology structure, we are able to extract the reasoning path $\mathcal{P}_{\text{valid}}$ that successfully connects from nodeRaw to nodeResult. Then, we effectively search the node that does not contribute to the final answer, which serves as a sign of redundancy in the reasoning process. Following the Eq. 8, we analyze the redundancy rate for each of the five LLMs on both the node and edge redundancy. Since the node and edge always show up together in our setting, their redundancy results are very similar; we just present redundancy on the ‘node’ level for conciseness in the left part of Figure. 3. We find that, surprisingly, the GPT-4o-mini shows a significantly high redundancy rate in both the nodes and edges, which might reveal that the high accuracy of the model comes from a broader searching space (or generating length) when conducting reasoning and proposing solutions. However, this also reflects that there would be great potential to optimize the reasoning process for the GPT-wise models. Comparatively, the DeepSeek (distilled llama version) is relatively low, which might indicate the model’s training was conducted with a special design to encourage the ‘valid’ reasoning, which eventually contributes to the final result. We show more results in the appendix for reference.
5.4 Findings on the Reasoning Patterns from LLMs (RQ3)
Based on the topology representations, we conducted further analysis in structural patterns, we use Graph2Vec (Narayanan et al., 2017) to embed the graph structure information, such as node connectivity, and perform K-means clustering on 4 classes and obtained silhouette score of 0.38, as shown in right part of Figure. 3, we found three major reasoning patterns of the LLMs (in color - red, green, purple on the right side of Figure. 3), they correspond with: ① Completion-based: the model has seen the data in the training phase, and completes it as a ‘cloze question’. For example, given the question that provides a feature of climate type, and asks for the answer as shown in Figure. 4, then in the sub-question step, the LLM directly asks ‘Does object A have such a feature?’, it shows a one-step, direct answer which possibly comes from the training memorization. ② Forward reasoning (actual reasoning): The model is actually performing reasoning based on the analysis procedure, like human beings, given a question, it first recalls relevant concepts and knowledge that contain the necessary information. Then, through a step-by-step narrowing-down process, it refines its understanding, gradually converging on the most precise answer. ③ Combination of Memory-based answering, forward reasoning (① + ②) ++ verify: The model uses two methods and double-checks to verify the answer. Three modes are shown in Figure. 4. We hope this could serve as an inspiration for the potential further study in answer topology.
Figure 4: The examples of three types of reasoning patterns. As shown in the image, the first type directly reflects the correct answer in its first reasoning step, and not actually performing the reasoning, the second type is like humans, try to recall the knowledge points, and narrow down the range, and find the possible answer in a thought chain, while the third one is a combination of two, it has the steps to ask for characteristics, but it also try to reflect if it has seen this question before at the edge 1, which is a direct attempt to use answer to match the question, the last step is to confirm if the memory of answer matches the requirement in the question.
6 Conclusion
In this paper, we introduce a novel framework Topo-UQ to quantify the uncertainty of LLMs’ reasoning through the explanations. By designing a reasoning topology elicitation module, we tackle the challenge of shaping the structural reasoning process of LLMs. Then, based on the constructed topologies, the paper provides a ‘graph edit distance’ based uncertainty measure named Reason-GED. It considers both the semantic meanings of the reasoning steps and the structural topology variance, by a set of well-designed operation costs: substitution and deletion (insertion), we can effectively quantify the reasoning uncertainty. The empirical studies show that this method provides better performance in revealing the true faithfulness of the natural language models. Besides, this paper also introduces a redundancy-based method to quantify the conciseness of the LLMs’ reasoning process, which could potentially serve as a guide for improving reasoning efficiency.
Limitations: Although the Topo-UQ shows an interesting direction (topology structure) to quantify the LLM reasoning uncertainty, it requires the LLMs to come with good instruction-following ability. Thus, this method is more applicable to larger models with long-context reasoning capabilities. Besides, the current method is validated by the faithfulness, but it does not concern the answer’s correctness from the reasoning path; analyzing the uncertainty relevant to the correctness would also be beneficial for real-world use-cases.
Ethics Statement
This paper focuses on the uncertainty quantification of the LLM explanations. The data adopted are the widely acknowledged public datasets, and contain no ethics sensitive information, which has no ethics concerns.
Reproducibility Statement
This project will release the codes for the topology elicitation that is used for uncertainty quantification, and we conducted pre-processing for the three datasets, which will also be updated to the repository: https://github.com/LongchaoDa/LLM-Topology.
@@ Appendix A Notation Table
Table 2: The notations and explanations in this paper.
Appendix B Reasoning Topology Construction Process in Detail
B.1 Knowledge Point Reflection Module
Template $T_{1}$: Given a question $x^{q}$}, reflect and generate the knowledge points as sub-questions necessary to solve this query. Ensure that the output is both sufficient and concise.
The model generates a set of knowledge points $K^{q}=M(x^{q},T_{1})=\{k_{1},k_{2},\dots,k_{n}\}$, where each $k_{i}$ corresponds to a specific sub-question or piece of information identified as necessary to address the query $x^{q}$ under the guidance of prompt $T_{1}$. To ensure traceability, we assign unique identifiers to each knowledge point using a tagging function $f(\cdot)$:
$K^{q}_{\text{tag}}=\{\text{id}_{1}:k_{1},\text{id}_{2}:k_{2},\dots,\text{id}_{n}:k_{n}\}.$ (9)
In the later sections of this paper, we assume the $k^{q}$ always carries its identifier while performing computing ($K^{q}\Leftrightarrow K^{q}_{tag}$).
B.2 Self-answering Module
$A=\{a_{1},a_{2},\dots,a_{n}\}=\{M\{k_{1},T_{2}\},M\{k_{2},T_{2}\},\dots,M\{k_{n},T_{2}\}\}$ (10)
$\mathcal{D}_{m}=\{(k_{1},a_{1}),(k_{2},a_{2}),\dots,(k_{n},a_{n})\}$ (11)
B.3 Reasoning Topology Construction Module
$\mathcal{\hat{D}}_{m}$ $=M(\mathcal{D}_{m},e^{F})=\{(a^{p_{1}},k_{1},a_{1}),(a^{p_{2}},k_{2},a_{2}),\dots,(a^{p_{n}},k_{n},a_{n})\}$ (12)
$(a^{p_{1}},k_{1},a_{1})\Rightarrow(\underbrace{a^{p_{1}},a_{1}}_{{nodes}},\underbrace{k_{1}}_{edge})$ (13)
where the order of two nodes is defined by the reasoning LLM.
Now, we can write a basic reasoning step as:
$\mathcal{\text{Step}}_{ij}=[\text{node}_{i},\text{edge}_{ij},\text{node}_{j}]$ (14)
$\mathcal{G}^{q}=(\mathcal{V},\mathcal{E}),$ (15)
where
$\mathcal{V}=\{\texttt{nodeRaw},\text{node}_{1},\dots,\texttt{nodeResult}\}=\{v_{0},v_{1},$...\} (16)
and the edges are expressed as:
$\mathcal{E}=\{e_{ij(1)},e_{ij(2)}$...\} (17)
where $e$ stands for edge and $\{e_{ij}\mid\text{edge}_{ij}:\text{node}_{i}\to\text{node}_{j}\}$, $e_{ij}$ represents reasoning operations or dependencies between nodes. (start with index ‘1’ since we assume ‘0’ is the nodeRaw). The graph-based structure captures the full reasoning topology, including branching, dependencies, and multi-step interactions, which allows for a better reflection of the relationships between intermediate steps.
Now, from a graph concept, the reasoning steps combined with Eq. 14 are formalized as below:
$S=\{\text{Step}_{ij}\mid\text{Step}_{ij}=[v_{i},e_{ij},v_{j}]v_{i},v_{j}\in\mathcal{V},e_{j}\in\mathcal{E}\}$ (18)
where each triplet represents a logical transition between reasoning steps. Note that for complex reasoning, the final answer does not necessarily rely on all of the reasoning steps. For example, when being asked about ”If it is currently summer in Australia, what season is it in Canada?” in the reasoning chain, some of the LLM might delve into ‘what causes the season differences, ’ which is a redundant step in concise reasoning.
Appendix C More Studies for Reasoning Graph Edit Distance
C.1 The case behind high $c_{\text{sub.}}$
Figure 5: The example of the cause of substitution cost: same sub-question, divergence answers.
C.2 The complete equation of deletion/insertion cost
Similarly to Eq. equation 5 for nodes, we define the deletion cost for an edge $e_{i}\in\mathcal{E}_{1}$ in graph $\mathcal{G}_{1}=(\mathcal{V}_{1},\mathcal{E}_{1})$ relative to the target graph $\mathcal{G}_{2}=(\mathcal{V}_{2},\mathcal{E}_{2})$ as:
$c_{\text{del}}(e_{i})=\frac{1}{2}\{\underbrace{\max_{e_{j}\in\mathcal{E}_{2}}\cos\big{(}\mathcal{L}(e_{i}),\mathcal{L}(e_{j})\big{)}}_{\text{Cross-graph Matching Term}}\;+\;\underbrace{1-\frac{1}{|\mathcal{E}_{1}|-1}\sum_{\begin{subarray}{c}e_{k}\in\mathcal{E}_{1}\\ e_{k}\neq e_{i}\end{subarray}}\cos\big{(}\mathcal{L}(e_{i}),\mathcal{L}(e_{k})\big{)}}_{\text{Internal Uniqueness Term}}\}.$ (19)
The Eq. equation 19 jointly considers the cross-graph matching cost and the internal uniqueness cost when deleting edge $e_{i}$.
  • Cross-graph Matching Term:
    $\max_{e_{j}\in\mathcal{E}_{2}}\cos\big{(}\mathcal{L}(e_{i}),\mathcal{L}(e_{j})\big{)}$
    measures the highest semantic similarity between the edge $e_{i}$ in $\mathcal{G}_{1}$ and any edge $e_{j}$ in the target graph $\mathcal{G}_{2}$. If $e_{i}$ has a very similar counterpart in $\mathcal{E}_{2}$, deleting it would remove important alignment information, leading to a higher deletion cost.
  • Internal Uniqueness Term:
    $1-\frac{1}{|\mathcal{E}_{1}|-1}\sum_{\begin{subarray}{c}e_{k}\in\mathcal{E}_{1}\\ e_{k}\neq e_{i}\end{subarray}}\cos\big{(}\mathcal{L}(e_{i}),\mathcal{L}(e_{k})\big{)}$
    calculates how unique edge $e_{i}$ is within its own graph $\mathcal{G}_{1}$. If $e_{i}$ is very similar (i.e., redundant) to many other edges, the average similarity will be high, making this term low; in other words, deleting a redundant edge is less costly.
Thus, the overall deletion cost for edge $e_{i}$ is the average of these two terms.
Appendix D Related Work
D.1 UQ for LLM
White-box Approaches
A significant body of research has focused on performing UQ for LLMs by inducing the models to output their uncertainty along with their responses Kadavath et al. (2022); Lin et al. (2022); Mielke et al. (2020); Tian et al. (2023). These methods often rely on token-level probabilities to train or fine-tune models for predicting uncertainty. While effective, these approaches require full access to the model’s structure and weights, which is impractical for black-box or commercial LLMs. For example, supervised methods such as those in Kadavath et al. (2022) estimate uncertainty using logits and ground truth labels but are computationally expensive and resource-intensive.
Black-box Approaches
Another line of work estimates uncertainty directly at the response level using semantic entropy Kuhn et al. (2023). While this method avoids token-level dependencies, it still relies on access to token probabilities, limiting its applicability in black-box settings. To address these limitations, researchers have proposed lightweight black-box methods that analyze response inconsistencies. For instance,  Lin et al. (2023) uses graph Laplacian eigenvalues as an uncertainty indicator, while  Chen & Mueller (2023) computes confidence scores from generated outputs to identify speculative or unreliable answers. However, these approaches primarily focus on semantic-level analysis and neglect the logical structure underlying reasoning processes. Moreover, methods like Lin et al. (2023) average entailment probabilities without considering directional information in reasoning paths.
Our work is agnostic to the white box or black box since it leverages the generated explanations as a proxy to measure the reasoning uncertainty as in Figure. 1. This enables a more nuanced and interpretable assessment of uncertainty in reasoning processes.
D.2 UQ for LLM Explanation
Explanation-based UQ focuses on assessing the reliability of natural language explanations (NLEs) generated by LLMs by either prompting models to express confidence in their explanations or analyzing consistency across multiple outputs under varying conditions Tanneru et al. (2024); Yadkori et al. (2024). While these methods provide insights into explanation robustness, they treat explanations to a question as unstructured text representation, which lacks structural information and fails to capture inconsistencies or leaps in logic. In contrast, our work explicitly leverages well-structured reasoning topologies to enhance the UQ process for explanations. This structured representation enables us to assess explanation uncertainties at a finer granularity within complex reasoning paths.
Appendix E Details of the GeoQA Dataset
@@
Appendix F Empirical Study on the Choice of Number for Generations
@@
Appendix G Details of Baseline methods
G.1 CoTA.
@@ Appendix G Details of Baseline methods
$\text{CoTA}(CoT_{a},CoT_{b})=$ $\ \frac{1}{N_{a}+N_{b}}\bigg{(}\sum_{i=1}^{N_{a}}\max_{j=1,\dots,N_{b}}E(s_{ai},s_{bj})$ (20)
@@
Table 3: Examples of GeoQA dataset.
@@
Glacier
These questions explore
glacial movement, erosion,
and the impact of climate
change on ice dynamics.
@@
Striations indicate past movement direction,
requiring cause-and-effect analysis,
while glacier mass loss demands understanding
the imbalance between accumulation and ablation.
@@
Earthquake
These questions focus on
seismic wave behavior,
fault activity, and
earthquake detection.
@@
Triangulation requires reasoning through
wave arrival times, while P-wave detection relies on
comparing wave speeds and impact
to explain early warning systems.
@@
permafrost
These questions examine
permafrost thawing,
climate feedback loops, and
seasonal variations.
@@
Thawing permafrost and methane release involve
feedback loops, while active layer variations
require analyzing environmental factors
like temperature and insulation.
@@
Monsoon
These questions cover
seasonal wind shifts,
monsoon patterns, and
storm formation.
@@
The East Asian monsoon’s impact on seasons
involves reasoning about wind shifts,
while monsoon depressions require
linking low-pressure systems to storm formation.
@@
G.4 NLI-logit-UQ.
NLI-logit-UQ is a variant of Entail-UQ that directly utilizes the raw logits from a Natural Language Inference (NLI) model to measure uncertainty, without applying the softmax operation. Given the set of explanations $\{a^{e}_{1},a^{e}_{2},\dots,a^{e}_{k}\}$, the NLI model computes a raw logit score for each pair of explanations:
$l_{ij}=\text{NLI}(a^{e}_{i},a^{e}_{j}),$
where $\text{NLI}(\cdot,\cdot)$ outputs a logit value representing the degree of entailment between $a^{e}_{i}$ and $a^{e}_{j}$.
In contrast to Entail-UQ, where these logits are passed through a softmax layer to obtain probabilities, the NLI-logit-UQ method uses the raw logits directly. The dissimilarity between explanations can be defined as:
$d_{ij}=1-l_{ij},$
or more generally, the uncertainty is computed as the variance of the raw logit values:
$\mathcal{U}(x^{q})=\text{Var}(\{l_{ij}\mid 1\leq i,j\leq k\}).$
Figure 6: The model’s success rate to generate legitimate reasoning topology. It is calculated by the percentage of LLMs successfully generating the reasoning path from the nodeRaw to nodeResult following the few-shot prompt. It can be witnessed that, generally the GeoQA is a harder one to generate due to the conditional question types, which are rarely seen in the LLM training tasks.
Appendix H Research Methods
H.1 Part One
@@
$\text{GED}(\mathcal{G}_{1},\mathcal{G}_{2})=C_{\text{sub.}}(\mathcal{P})+C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})$ (21)
@@
$C_{\text{sub.}}(\mathcal{P})=\sum_{(v_{i},v_{j})\in\mathcal{P}_{v}}c(v_{i},v_{j})+\sum_{(e_{i},e_{j})\in\mathcal{P}_{e}}c(e_{i},e_{j}),$ (22)
where $c(v_{i},v_{j})$ and $c(e_{i},e_{j})$ represent the node and edge substitution costs, respectively. The term $C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})$ captures the deletion costs for nodes and edges in $\mathcal{G}_{1}$ that are not matched, given by:
$C_{\text{del.}}({\mathcal{P}})=C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})=\sum_{v_{i}\in\mathcal{V}_{1}\setminus\mathcal{P}_{v}}c_{\text{del.}}(v_{i})+\sum_{e_{i}\in\mathcal{E}_{1}\setminus\mathcal{P}_{e}}c_{\text{del.}}(e_{i}).$ (23)
Here, $c_{\text{del.}}(v_{i})$ and $c_{\text{del.}}(e_{i})$ denote the deletion costs of unmatched nodes and edges, respectively. This formulation quantifies the total cost required to align the two reasoning structures by summing the substitution costs of matched components and the deletion costs of unmatched ones.
H.2 Algorithm for Detection of Redundant Nodes and Dead Branches
@@
$\text{Out}(v_{k})=\emptyset\quad\text{and}\quad v_{k}\neq\text{NodeResult}.$ (24)
@@
$v_{k}\notin\bigcup_{[v_{i},v_{j},e_{ij}]\in\mathcal{P}_{\text{valid}}}\{v_{i},v_{j}\}.$ (25)
@@
Appendix I Experimental Details
I.1 Details of the dataset
@@
  • GeoQA: GeoQA is a self-constructed dataset designed to evaluate the reasoning capabilities of LLMs in conditional questions. With 20 categories, including climate and tectonic processes, its tasks require inference of specific results from given conditions. GeoQA emphasizes multi-step reasoning and domain-specific integration, making it a challenging benchmark (details in the Appendix E).
I.2 The calculation of the faithfulness core